OpenStudy (anonymous):

cosxsinx=sin^2x, solve for x.

OpenStudy (anonymous):

\[cosxsinx= \sin^2x\], solve for x!

OpenStudy (phi):

if you had xy = y^2 you would write it as y^2 -x y= 0- y( y-x)=0 and y=0 or y=x are the two answers can you use this idea to solve your problem?

OpenStudy (anonymous):

cosxsinx-sin^2x=0 sinx(cosx-sinx)=0 sinx=0 cosx-sinx=0 x=pi, 0 x= ?

OpenStudy (phi):

for the second equatin divide by cos x 1- sin x/ cos x =0 1 - tan x =0 tan x =1 so all angles that make tan x =1